# Geodesically complete universes

### v1

@@ I Introduction
However, the current consensus is that some of these assumptions, and most likely the very framework provided by GR, fails to describe adequately the very early universe, at least because quantum gravitational effects are expected to become relevant before reaching a singularity. Understanding the precise form in which this happens is subject of intense research, with the development of diverse theoretical models [8, 9, 10, 11, 12, 13, 14, 15, 16] that could eventually be tested observationally [17, 18, 19, 20].
@@ II Beyond Penrose’s theorem
The theorem then proceeds in two steps. The first one consists in showing that an initial negative expansion and the null convergence condition are enough to prove, via the Raychaudhuri equation, that the expansion will become infinitely negative in the future for some finite value of the affine parameter of a null congruence orthogonal to ${\mathcal{T}}^{2}$. This focusing point is expected to be singular. The second step consisted in a ingenious use of topological arguments to prove that assuming the regularity of the focusing point and the consequent compactness of the boundary of the chronological future of ${\mathcal{T}}^{2}$ is incompatible with the presence of a non-compact Cauchy hypersurface.
@@ III Trapped regions in FLRW geometries
Figure 2 shows the causal structure of a decelerating FLRW metric and its trapped region, which will exist for the open case only for $\dot{a}>1$. Trapped surfaces appear for sufficiently large distances to the reference point $r=0$ being used. However, due to homogeneity and isotropy, the choice of reference point is fiduciary and has no physical meaning. This is different with respect to spherically-symmetric black holes, in which the gravitational potential has a defined center. Hence, in a FLRW spacetime it is enough to show that a point belongs to a trapped surface to conclude that all points belong to trapped surfaces. In other words, there are no trapped regions in the usual sense: either all points are trapped, or no points are trapped.
@@
As a consequence, $a(t)$ must be at least a $C^{2}$ function to avoid curvature singularities.
@@ V Regular Flat FLRW universes
Hence, spacetimes in this class are geodesically complete if and only if the derivative of the scale factor vanishes at some $t=t_{0}$, while the scale factor remains non-zero at $t=t_{0}$. Indeed, the limiting case in which also the scale factor vanishes is not admissible — as it would entail a manifest curvature singularity. For $t<t_{0}$, in order to avoid the formation of further trapped surfaces, $\dot{a}(t)\leqslant 0$. Hence, spacetimes in this class describe a bounce between a contracting and an expanding universe where a future trapped region is continuously connected with a past trapped region, being both delimited by the trapping horizon $r=\pm{1}/{a(t)}$.
@@
Note that we are implicitly assuming that the scale factor is an analytic function of time which, if dropped, would allow also the possibility that $a(t)$ vanishes in a closed interval or half-line (in the latter case, the geometry describes an expanding universe emerging from a stationary phase [8]). Analytical solutions that are effectively emergent are however possible, for example the expansion can be preceded by an oscillatory phase in which $\dot{a}$ is 0 on average (see [11] for a specific realization). These solutions embody a hybrid nature, drawing from both emergent and bouncing universe scenarios, as $\dot{a}$ assumes also negative values during the “emergent” phase.
@@
Example: Let us require to have a regular expanding universe at any time thus $\dot{a}(t)\geqslant 0$ and $a(t)>0$ for any $t$. In order to have the defocusing at infinite affine distance we then need $\lim_{t\rightarrow-\infty}\dot{a}(t)=0$. A simple analytic profile for $a(t)$ is then $a(t)=a_{0}+b_{0}e^{Ht}$, that is a sort of “regularized inflation”.
@@
where we introduced $g(t_{\rm B2})=r(t_{\rm B2})+{t_{\rm B2}}/{a_{-\infty}}+b_{1}\ln{|t_{\rm B2}|}-b_{2}/{t_{\rm B2}}$ is not a function of $t$. Since the integrand near goes to 0 in the $t\rightarrow-\infty$ limit at most as ${1}/{t}$, the integral in the last line is divergent.
@@ VII.3 Case C
Figure 11: Scheme of the allowed closed homogeneous and isotropic regular geometries. These geometries as follows: (i) the value of the affine parameter at which all outgoing geodesics are defocused $\lambda_{def}$, so that $\theta_{+}(\lambda_{def})=0$; (ii) $a(\lambda_{def})$ or equivalently the value of both expansions for the same value of the affine parameter [note that, if $a(\lambda_{def})=0$, then $\theta_{\pm}(\lambda_{def})=-\infty$]; (iii) the derivative of the scale factor for $\lambda>\lambda_{def}$, i.e. for $t<t_{def}$.
@@ VIII Conclusions
The analysis of closed universes presented unique challenges due to the violation of the non-compactness assumption and to the diverging nature of certain functions influencing geodesic expansions. Despite this, it was established that, also in this case, true defocusing points at finite or infinite affine distance are contingent upon specific behaviors of the scale factor derivative with respect to time. Moreover, the absence of defocusing leads to the inevitable occurrence of a curvature singularity.
Regarding the physical interpretation of the bouncing solutions, it is interesting to keep in mind the following consideration. The metric alone does not provide any real information regarding the direction in which time flows. GR is time-symmetric, as flipping the $t$-direction on a globally hyperbolic manifold results in another valid solution to the field equations. [^4] Note that, due to the fact that at the bounce $\dot{a}(t)=0$, this would be possible while keeping the metric components at least $C^{1}$, and could be possibly fine-tuned so that at the bounce also has $\ddot{a}(t)=0$ so as to ensure regularity of curvature tensors as well.

### v2

@@ I Introduction
However, the current consensus is that some of these assumptions, and most likely the very framework provided by GR, fails to describe adequately the very early universe, at least because quantum gravitational effects are expected to become relevant before reaching a singularity. Understanding the precise form in which this happens is subject of intense research, with the development of diverse theoretical models [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] that could eventually be tested observationally [22, 23, 24, 25].
@@ II Beyond Penrose’s theorem
The theorem then proceeds in two steps. The first one consists in showing that an initial negative expansion and the null convergence condition are enough to prove, via the Raychaudhuri equation, that the expansion will become infinitely negative in the future for some finite value of the affine parameter of a null congruence orthogonal to ${\mathcal{T}}^{2}$. This focusing point is expected to be singular. The second step consisted in a ingenious use of topological arguments to prove that assuming the regularity of the focusing point is incompatible with the presence of a non-compact Cauchy hypersurface.
@@
In this work we analyze and classify globally hyperbolic and time-orientable manifolds describing expanding, homogeneous and isotropic regular universes. In order to avoid the formation of a singularity, at least one of the assumptions of the Penrose theorem must necessarily be violated. For flat and open manifolds the last two assumptions of the theorem will always be verified and thus the geodesically complete geometries that we will find will necessarily violate the null convergence condition. For closed manifolds similar considerations imply that all the geodesically complete geometries that we find necessarily violate the timelike convergence condition (see section VII).
@@ III Trapped regions in FLRW geometries
Fig. 2 shows the causal structure of a decelerating FLRW metric and its trapped region, which will exist for the open case only for $\dot{a}>1$. Trapped surfaces appear for sufficiently large distances to the reference point $r=0$ being used. However, due to homogeneity and isotropy, the choice of reference point is fiduciary and has no physical meaning. This is different with respect to spherically-symmetric black holes, in which the gravitational potential has a defined center. Hence, in a FLRW spacetime it is enough to show that a point belongs to a trapped surface to conclude that all points belong to trapped surfaces. In other words, there are no trapped regions in the usual sense: either all points are trapped, or no points are trapped.
@@
As a consequence, a necessary condition to avoid curvature singularities is for $a(t)$ to be at least a $C^{2}$ function.
@@ V Regular Flat FLRW universes
Hence, spacetimes in this class are geodesically complete if and only if the derivative of the scale factor vanishes at some $t=t_{0}$, while the scale factor remains non-zero at $t=t_{0}$. Indeed, the limiting case in which also the scale factor vanishes is not admissible — as it would entail a manifest curvature singularity. For $t<t_{0}$, in order to avoid the formation of further trapped surfaces, $\dot{a}(t)\leqslant 0$. Hence, spacetimes in this class describe a bounce between a contracting and an expanding universe (Fig. 3) where a future trapped region is continuously connected with a past trapped region, being both delimited by the trapping horizon $r=\pm{1}/{a(t)}$.
@@
Note that we are implicitly assuming that the scale factor is an analytic function of time which, if dropped, would allow also the possibility that $a(t)$ vanishes in a closed interval or half-line (in the latter case, the geometry describes an expanding universe emerging from a stationary phase [8]). Analytical solutions that are effectively emergent are however possible, for example the expansion can be preceded by an oscillatory phase in which $\dot{a}$ is 0 on average (see [11] for a specific realization). These solutions embody a hybrid nature, drawing from both emergent and bouncing universe scenarios, as $\dot{a}$ assumes also negative values during the “emergent” phase (Fig. 4).
The causal structure of the spacetimes entering this class is shown in Fig. 5.
@@
Example: Let us require to have a regular expanding universe at any time thus $\dot{a}(t)\geqslant 0$ and $a(t)>0$ for any $t$. In order to have the defocusing at infinite affine distance we then need $\lim_{t\rightarrow-\infty}\dot{a}(t)=0$. A simple analytic profile for $a(t)$ is then $a(t)=a_{0}+b_{0}e^{Ht}$, that is a sort of “regularized inflation” (see Fig. 6 and Fig. 7).
@@
where we introduced $g(t_{\rm B2})=r(t_{\rm B2})+{t_{\rm B2}}/{a_{-\infty}}+b_{1}\ln{|t_{\rm B2}|}-b_{2}/{t_{\rm B2}}$ that is not a function of $t$. Since the integrand goes to 0 in the $t\rightarrow-\infty$ limit at most as ${1}/{t}$, the integral in the last line is divergent.
@@ VII.3 Case C
Figure 11: Scheme of the allowed closed homogeneous and isotropic regular geometries. These geometries are classified according to: (i) the value of the affine parameter at which all outgoing geodesics are defocused $\lambda_{def}$, so that $\theta_{+}(\lambda_{def})=0$; (ii) $a(\lambda_{def})$ or equivalently the value of both expansions for the same value of the affine parameter [note that, if $a(\lambda_{def})=0$, then $\theta_{\pm}(\lambda_{def})=-\infty$]; (iii) the derivative of the scale factor for $\lambda>\lambda_{def}$, i.e. for $t<t_{def}$.
@@ VIII Conclusions
Some concrete examples for these three scenarios are constructed and studied in [19] where it is explicitly shown that they all violate the null energy conditions for at least some amount of time.
The analysis of closed universes presented unique challenges due to the violation of the non-compactness assumption and to the diverging nature of certain functions influencing geodesic expansions. Despite this, it was established that, also in this case, true defocusing points at finite or infinite affine distance are contingent upon specific behaviors of the scale factor derivative with respect to time. Moreover, the absence of defocusing leads to the inevitable occurrence of a curvature singularity. Regarding the physical interpretation of the bouncing solutions, it is interesting to keep in mind the following consideration. The metric alone does not provide any real information regarding the direction in which time flows. GR is time-symmetric, as flipping the $t$-direction on a globally hyperbolic manifold results in another valid solution to the field equations. [^5] Note that, due to the fact that at the bounce $\dot{a}(t)=0$, this would be possible while keeping the metric components at least $C^{1}$, and could be possibly fine-tuned so that at the bounce also has $\ddot{a}(t)=0$ so as to ensure regularity of curvature tensors as well.